# Adaptive Federated Learning in Heterogeneous Wireless Networks with Independent Sampling

### v2

@@ II-A Unbiased FL with Independent Client Sampling
2       At each client $n=1,2,...,N$, do:  Sample $\mathbb{I}_{r}^{n}\sim\text{Bernoulli}(q_{n})$;   // Sampling  Update gradients;      // Computation  Send $-\frac{\eta\mathbb{I}_{r}^{n}}{q_{n}}g_{n}(x_{r})$ to server;  // Communication At the server, do:  Update $\!x_{r+1}\!\leftarrow\!x_{r}\!-\!\sum_{n=1}^{N}\!a_{n}\frac{\eta\mathbb{I}_{r}^{n}}{q_{n}}g_{n}(\!x_{r}\!)$; // Aggregation  Send $x_{r+1}$ to all clients;   // Synchronization
@@ IV-C Optimization Algorithm for q
1 Server runs Algorithm 1 with uniform sampling $\textbf{q}_{1}$ and full sampling $\textbf{q}_{2}$, respectively; Server recored $R_{1}$ and $R_{2}$, when reaching $F_{s}$; Calculate $\alpha$ and $\beta$ using (14); for $M(\epsilon)=M_{min},M_{min}+\epsilon,M_{min}+2\epsilon,...,M_{max}$ do
2       Substitute $N,\tau_{n},t_{n},f_{tot},a_{n},M(\epsilon),\alpha,\beta$ into P3; Solve P3 via CVX, and obtain $\textbf{q}^{*}(M(\epsilon))$;
@@ V-C Results
Figure 2: CNN@MNIST
Figure 3: MobileNet@CIFAR10
Figure 4: LSTM@Shakespeare

### v3

@@ II-A Unbiased FL with Independent Client Sampling
2       At each client $n=1,2,...,N$, do:
3       Sample $\mathbb{I}_{r}^{n}\sim\text{Bernoulli}(q_{n})$;   // Sampling
4       Update gradients;      // Computation
5       Send $-\frac{\eta\mathbb{I}_{r}^{n}}{q_{n}}g_{n}(x_{r})$ to server;  // Communication
6      At the server, do:
7       Update $\!x_{r+1}\!\leftarrow\!x_{r}\!-\!\sum_{n=1}^{N}\!a_{n}\frac{\eta\mathbb{I}_{r}^{n}}{q_{n}}g_{n}(\!x_{r}\!)$; // Aggregation
8       Send $x_{r+1}$ to all clients;   // Synchronization
@@ IV-C Optimization Algorithm for q
1
2Server runs Algorithm 1 with uniform sampling $\textbf{q}_{1}$ and full sampling $\textbf{q}_{2}$, respectively;
3Server recored $R_{1}$ and $R_{2}$, when reaching $F_{s}$;
4Calculate $\alpha$ and $\beta$ using (14);
5for $M(\epsilon)=M_{min},M_{min}+\epsilon,M_{min}+2\epsilon,...,M_{max}$ do
6       Substitute $N,\tau_{n},t_{n},f_{tot},a_{n},M(\epsilon),\alpha,\beta$ into P3;
7      Solve P3 via CVX, and obtain $\textbf{q}^{*}(M(\epsilon))$;